In a geometric progression, the first, third and fifth terms may be considered as the first, fourth, and sixteenth terms of an arithmetic progression. Determine the fourth term of this arithmetic progression knowing that its first term is 5.

Grade:11

1 Answers

Vikas TU
7 years ago
a = 5 given.
and 
ar^2 = a + 3d...............(1)
and
ar^4 = a + 15d.................(2)
solve eqns. 1 and 2 after substituting a=5
and write T4 = a+3d.